# An Efficient Intelligent Semi-Automated Warehouse Inventory Stocktaking System

### v2

@@ Abstract
As supply chain management continues to evolve, efficient inventory management systems have become increasingly crucial. However, traditional manual methods often struggle to meet the complexities of modern market demands, particularly when it comes to data accuracy, delays in monitoring, and the heavy reliance on subjective experience for forecasting. This study introduces an intelligent, semi-automated barcode-based inventory management system designed to overcome these challenges. The system integrates barcode technology with a distributed architecture, combined with big data analytics and machine learning for real-time tracking and accurate inventory predictions Its performance has been validated through multiple simulation tests, where it has outperformed traditional RFID technology in certain cases. Through careful system design, technology exploration, and validation, this research demonstrates the significant potential of this intelligent system in improving inventory management efficiency and accuracy.
Keywords stocktaking $\cdot$ inventory  $\cdot$ stock count  $\cdot$ RFID  $\cdot$ semi-automated  $\cdot$ warehouse
@@ 1 Introduction
In recent years, fully automated inventory technologies have experienced significant advancements, with the utilization of next-generation ultra-high-frequency RFID (Radio-Frequency Identification) employing radio waves for real-time identification and tracking of tagged objects. The application of RFID in warehouse management traces back to the 1990s, with initiatives by major players such as Walmart and the US Department of Defense. However, it wasn’t until the 2000s that RFID technology gained widespread adoption due to the establishment of standards and cost reductions [13]. While new RFID systems have demonstrated impressive accuracy in real-time inventory monitoring and complex tasks like expiration date tracking, the full extent of their potential remains unpredictable. This raises questions about the limited adoption of these powerful systems in large manufacturing factories.
Although fully automated inventory technology is promising, warehouse management is not yet ready for full adoption.. Prior research, such as Dong et al. [20], has highlighted a fundamental issue in large-scale RFID deployments: the complexity of load balancing in dense inventory environments. Their analysis demonstrated that minimizing RFID reader overload is an NP-hard problem, meaning that maintaining consistent performance in high-density warehouses is computationally intensive and inherently inefficient. This challenge is further exacerbated by environmental factors—Griffiths et al. [8] found that RFID accuracy significantly deteriorates when tags are placed near metals or liquids, making RFID unreliable in diverse warehouse conditions.
Beyond technical limitations, RFID adoption also faces economic and operational barriers. Huber et al. [7] identified that the high cost of RFID tags, readers, and integration with existing warehouse management systems creates a financial burden for businesses. While RFID promises real-time stock tracking, achieving full automation often requires substantial infrastructure investment, which many companies find prohibitive. Furthermore, studies such as White et al. [15] indicate that fully automated RFID inventory systems may not be as cost-effective as initially anticipated, due to ongoing maintenance expenses and frequent calibration requirements.
@@
By critically assessing RFID’s limitations and offering a structured alternative, this research provides a clear pathway for businesses seeking to modernize their inventory management without the financial and operational burdens associated with full RFID adoption. This study contributes to the field by presenting a scalable and pragmatic inventory solution, offering valuable insights for warehouse managers and supply chain professionals navigating the transition toward digitalized stock control.
2 Related Work and Research Gap
Barcode-based inventory management systems have long been a staple in warehouse operations due to their affordability and simplicity. However, despite their widespread use, they exhibit several limitations, especially in large-scale or high-turnover environments. One major issue with traditional barcode systems is the reliance on batch synchronization for inventory updates. As a result, data is often outdated by the time it is processed, which hinders timely decision-making. The manual scanning process, though essential for the accuracy of the data, introduces delays as updates occur only once the scanning process is complete.
The dependence on human labor also opens the door to various types of errors. Barcode scanning systems are still heavily reliant on operators manually scanning each item, which leads to mistakes such as missed scans, misidentification of items, or duplication of scans. Such errors accumulate over time, significantly affecting inventory accuracy and requiring additional manual checks to resolve discrepancies. As warehouses grow, these inefficiencies become more pronounced, limiting the scalability and operational effectiveness of barcode-based systems.
@@ 2.1 Semi-Automated Barcode-Based Systems
Semi-automated barcode-based systems, while still reliant on manual scanning, have been increasingly integrated with software systems that streamline the inventory process. Research has shown that integrating barcode scanning with cloud-based platforms can help improve data accuracy and enable real-time synchronization between the warehouse and ERP systems.An study addresses the inventory management issues of the Zambia Air Force (ZAF) by proposing an automated inventory management system based on cloud architecture and barcode technology to resolve problems caused by manual operations, such as inventory errors, inefficient management, item losses, and high costs [21]. System testing demonstrated that barcode scanning improved inventory accuracy by 95%, was over three times faster than manual management, and enhanced security through CCTV monitoring. However, the study has certain limitations, including the lack of consideration for more advanced technologies such as RFID, insufficient security analysis, unquantified CCTV effectiveness, and the absence of a return on investment (ROI) assessment.
2.2 RFID-Based Semi-Automated Inventory Counting
@@
Similarly, RFID represents the latest technology in next-generation warehouse inventory management and has been included in comparative studies to evaluate its advantages and limitations in improving inventory management efficiency. A study [15] conducted an experimental comparison between barcoding and radio frequency identification (RFID) technologies in inventory management, focusing on scanning speed, accuracy, equipment failure rate, and applicable environments, with testing carried out in a cold chain warehouse. The results showed that RFID scanning speed was 2.5 times faster than barcoding, but its error rate (46.5%) and equipment failure rate (37%) were higher, and its performance in cold chain environments was unstable. The study also highlighted that two-dimensional barcodes (2D Barcode) are narrowing the performance gap with RFID and suggested that future research should explore hybrid systems (Hybrid Systems) that integrate the advantages of both technologies. However, the study lacked detailed cost analysis and long-term return on investment (ROI) evaluation and was limited to cold chain warehouses, without addressing the needs of other industries. Future advancements in expanding the scope of experiments, optimizing RFID equipment, researching hybrid systems, and analyzing its potential in supply chain big data management could enhance the intelligence and efficiency of inventory management. However, while these solutions address some of the shortcomings of traditional barcode systems, they still rely heavily on manual scanning and lack intelligent anomaly detection capabilities, leaving them vulnerable to human error and delayed issue resolution.
@@
In session 2.21, Dong’s research revealed that load balancing problem in large-scale RFID systems reveals that RFID readers in high-density tag environments are prone to uneven load distribution, data loss, and excessive energy consumption. However, the study does not explore dynamic environments, cost analysis, and security issues in depth, which share certain similarities with the conclusions drawn by White et al. (2007) in their comparison of barcoding and RFID technologies [15], which also served as one of the motivations for this research.
@@
  • RFID Tags: Embedded microchips attached to inventory items for unique identification.
  • RFID Readers: Devices that emit radio waves to activate and read RFID tags.
@@
    Readability Issues in Complex Warehouse Environments
@@
    $P_{\text{read}}=\begin{cases}1-\frac{N_{\text{tags}}}{N_{\text{max}}},&\text{if }N_{\text{tags}}\leq N_{\text{max}},\\ 0,&\text{if }N_{\text{tags}}>N_{\text{max}}.\end{cases}$ (1)
@@
    $P_{\text{loss}}=1-\frac{C_{\text{max}}}{M\times R}$ (2)
    where $C_{\text{max}}$ is server processing capacity, $M$ is the number of RFID readers, and $R$ is the data rate per reader.
@@
    Unlike barcode-based solutions, RFID requires specialized ERP connectors for real-time data processing. Many legacy warehouse management systems lack native RFID support, necessitating additional middleware.
@@ 2.3 Research Gap
To fill this gap, our system combines the best of barcode technology with cutting-edge stream processing and AI capabilities, offering a more balanced solution that minimizes costs, reduces errors, and scales efficiently.
We compare our system to RFID-based, traditional manual inventory counting, and barcode-based systems in the following table:
@@
Feature RFID-Based System Traditional Manual Stocktaking Proposed System (Flutter + Flink + AI)
@@
Stocktaking Speed Fast (but affected by signal interference) Slow (Manual item-by-item scanning) Fast (Automated barcode scanning + AI anomaly detection)

### v3

# Chunan Tong, CSCP-F

@@ Abstract
As supply chain management continues to evolve, efficient inventory management systems have become increasingly crucial. However, traditional manual methods often struggle to meet the complexities of modern market demands, particularly when it comes to data accuracy, delays in monitoring, and the heavy reliance on subjective experience for forecasting. This study introduces an intelligent, semi-automated barcode-based inventory management system designed to overcome these challenges. The system integrates barcode technology with a distributed architecture, combined with big data analytics and machine learning for real-time tracking and accurate inventory predictions. Its performance has been validated through multiple simulation tests, where it has outperformed traditional RFID technology in certain cases. Through careful system design, technology exploration, and validation, this research demonstrates the significant potential of this intelligent system in improving inventory management efficiency and accuracy.
keywords:
stocktaking, inventory, stock count, RFID, semi-automated, warehouse
@@ 1 INTRODUCTION
In recent years, fully automated inventory technologies have experienced significant advancements, with the utilization of next-generation ultra-high-frequency RFID (Radio-Frequency Identification) employing radio waves for real-time identification and tracking of tagged objects. The application of RFID in warehouse management traces back to the 1990s, with initiatives by major players such as Walmart and the US Department of Defense. However, it wasn’t until the 2000s that RFID technology gained widespread adoption due to the establishment of standards and cost reductions [1]. While new RFID systems have demonstrated impressive accuracy in real-time inventory monitoring and complex tasks like expiration date tracking, the full extent of their potential remains unpredictable in research of [17]. This raises questions about the limited adoption of these powerful systems in large manufacturing factories.
Although fully automated inventory technology is promising, warehouse management is not yet ready for full adoption. Prior research, such as [19], has highlighted a fundamental issue in large-scale RFID deployments: the complexity of load balancing in dense inventory environments. Their analysis demonstrated that minimizing RFID reader overload is an NP-hard problem, meaning that maintaining consistent performance in high-density warehouses is computationally intensive and inherently inefficient. This challenge is further exacerbated by environmental factors—[8] found that RFID accuracy significantly deteriorates when tags are placed near metals or liquids, making RFID unreliable in diverse warehouse conditions.
Beyond technical limitations, RFID adoption also faces economic and operational barriers. [7] identified that the high cost of RFID tags, readers, and integration with existing warehouse management systems creates a financial burden for businesses. While RFID promises real-time stock tracking, achieving full automation often requires substantial infrastructure investment, which many companies find prohibitive. Furthermore, studies such as [14] indicate that fully automated RFID inventory systems may not be as cost-effective as initially anticipated, due to ongoing maintenance expenses and frequent calibration requirements.
@@
By critically assessing RFID’s limitations and offering a structured alternative, this research provides a clear pathway for businesses seeking to modernize their inventory management without the financial and operational burdens associated with full RFID adoption. This study contributes to the field by presenting a scalable and pragmatic inventory solution, offering valuable insights for warehouse managers and supply chain professionals navigating the transition toward digitalized stock control [27].
2 LITERATURE REVIEW
Barcode-based inventory management systems have long been a staple in warehouse operations due to their affordability and simplicity in [13]. However, despite their widespread use, they exhibit several limitations, especially in large-scale or high-turnover environments. One major issue with traditional barcode systems is the reliance on batch synchronization for inventory updates. As a result, data is often outdated by the time it is processed, which hinders timely decision-making. The manual scanning process, though essential for the accuracy of the data, introduces delays as updates occur only once the scanning process is complete.
The dependence on human labor also opens the door to various types of errors. Barcode scanning systems are still heavily reliant on operators manually scanning each item, which leads to mistakes such as missed scans, misidentification of items, or duplication of scans. Such errors accumulate over time, significantly affecting inventory accuracy and requiring additional manual checks to resolve discrepancies in [9]. As warehouses grow, these inefficiencies become more pronounced, limiting the scalability and operational effectiveness of barcode-based systems.
@@ 2.1 Semi-Automated Barcode-Based Systems
Semi-automated barcode-based systems, while still reliant on manual scanning, have been increasingly integrated with software systems that streamline the inventory process. Research has shown that integrating barcode scanning with cloud-based platforms can help improve data accuracy and enable real-time synchronization between the warehouse and ERP systems. A study by [20] addresses the inventory management issues of the Zambia Air Force (ZAF) by proposing an automated inventory management system based on cloud architecture and barcode technology to resolve problems caused by manual operations, such as inventory errors, inefficient management, item losses, and high costs. System testing demonstrated that barcode scanning improved inventory accuracy by 95%, was over three times faster than manual management, and enhanced security through CCTV monitoring. However, the study has certain limitations, including the lack of consideration for more advanced technologies such as RFID, insufficient security analysis, unquantified CCTV effectiveness, and the absence of a return on investment (ROI) assessment.
2.2 RFID-Based Semi-Automated System
@@
Similarly, RFID represents the latest technology in next-generation warehouse inventory management and has been included in comparative studies to evaluate its advantages and limitations in improving inventory management efficiency. A study by [14] conducted an experimental comparison between barcoding and radio frequency identification (RFID) technologies in inventory management, focusing on scanning speed, accuracy, equipment failure rate, and applicable environments, with testing carried out in a cold chain warehouse. The results showed that RFID scanning speed was 2.5 times faster than barcoding, but its error rate (46.5%) and equipment failure rate (37%) were higher, and its performance in cold chain environments was unstable. The study also highlighted that two-dimensional barcodes (2D Barcode) are narrowing the performance gap with RFID and suggested that future research should explore hybrid systems that integrate the advantages of both technologies. However, the study lacked detailed cost analysis and long-term return on investment (ROI) evaluation and was limited to cold chain warehouses, without addressing the needs of other industries. Future advancements in expanding the scope of experiments, optimizing RFID equipment, researching hybrid systems, and analyzing its potential in supply chain big data management could enhance the intelligence and efficiency of inventory management. However, while these solutions address some of the shortcomings of traditional barcode systems, they still rely heavily on manual scanning and lack intelligent anomaly detection capabilities, leaving them vulnerable to human error and delayed issue resolution.
@@
[19] revealed that the load balancing problem in large-scale RFID systems shows that RFID readers in high-density tag environments are prone to uneven load distribution, data loss, and excessive energy consumption. However, the study does not explore dynamic environments, cost analysis, and security issues in depth, which share certain similarities with the conclusions drawn by [14], which also served as one of the motivations for this research.
@@
  • RFID Tags: embedded microchips attached to inventory items for unique identification,
  • RFID Readers: devices that emit radio waves to activate and read RFID tags, and
@@
This technology sometimes encounter several challenges, including:
@@
    Readability Issues in Complex Environments
@@
    $P_{\text{read}}=\begin{cases}1-\frac{N_{\text{tags}}}{N_{\text{max}}},&\text{if }N_{\text{tags}}\leq N_{\text{max}},\\ 0,&\text{if }N_{\text{tags}}>N_{\text{max}},\end{cases}$ (1)
@@
    $P_{\text{loss}}=1-\frac{C_{\text{max}}}{M\times R},$ (2)
    where $C_{\text{max}}$ is server capacity, $M$ is the number of RFID readers, and $R$ is the data rate per reader.
@@
    Unlike barcode-based solutions, RFID requires specialized ERP connectors for real-time processing. Many legacy warehouse management systems lack native RFID support, necessitating additional middleware.
@@
2.3 Cutting-Edge Technologies in Stocktaking
2.3.1 IoT-Based Inventory Tracking System
The Internet of Things (IoT) represents a transformative approach to real-time inventory management by enabling the interconnection of physical assets through wireless networks, sensors, and cloud-based platforms. IoT systems integrate various components such as RFID, BLE, GPS, LPWAN (LoRa, NB-IoT, LTE-M), and smart devices to collect, process, and transmit inventory-related data [21].
Table 1: Comparison of LoRa, NB-IoT, and LTE-M in Inventory Auditing
Technology Frequency Band and Power Consumption Coverage Range Data Rate and Latency Applications
LoRa (LoRaWAN) Unlicensed spectrum, ultra-low power (1 mAh to join; 44-byte uplink: 100 $\mu$Ah) Up to 20 km rural, 5 km urban Low rate (up to 50 kbps); high latency (2–3.5 sec) Static inventory tracking (e.g., fixed assets in warehouses)1
NB-IoT Licensed LTE spectrum, moderate power (3 mAh to join; 44-byte uplink: 1.8 mAh) $\sim$4 km rural, 1 km urban Higher throughput (up to 200 kbps); low latency ($\sim$244 ms) Dynamic inventory tracking (e.g., shelf monitoring in retail)2
LTE-M Licensed spectrum, higher power; supports mobility Similar to NB-IoT; optimized for mobility Higher data rate (up to 200 kbps); lower latency Mobile inventory tracking (e.g., goods-in-transit)3
1Semtech used LoRaWAN for airport baggage tag tracking, reducing costs by 30–40%.
2Link Labs demonstrated reliable NB-IoT applications in warehouse asset tracking.
3Widely used in asset tracking with high update frequency and data rate requirements.
Among the enabling technologies, LPWAN solutions such as LoRa, NB-IoT, and LTE-M are particularly suited for large-scale inventory tracking due to their long-range connectivity and low power consumption. Complementary technologies like RFID and BLE also play vital roles: RFID excels in short-range, high-precision tracking (e.g., shelf management in warehouses), while BLE supports low-power indoor positioning (e.g., retail asset monitoring). For outdoor scenarios, GPS is often paired with LPWAN to track mobile inventory, such as goods in transit. Table 1 compares the technical characteristics and applications of LoRa, NB-IoT, and LTE-M in inventory auditing.
In sum, the effectiveness of IoT technologies in the market depends on factors such as the distance from the base station and signal density, similar to mobile phones [24]. In many cases, warehouses are situated in enclosed environments, which can lead to poor signal reception, hindering IoT devices from transmitting their heartbeat packets back to the server [25, 27]. Additionally, while advanced LoRa IoT nodes can communicate with each other, this inter-node communication may cause network congestion in both uplink and downlink channels, potentially leading to system failures [27]. However, a significant advantage of IoT devices is their low power consumption; they often send heartbeat packets infrequently, allowing batteries to last over five years [26].
2.3.2 Inventory Tracking System Using Drone
Drone-assisted inventory systems, leveraging Unmanned Aerial Vehicles (UAVs), represent a transformative approach to automated warehouse management tasks such as stocktaking and item tracking. These systems are particularly advantageous in large-scale facilities. The primary applications include:
  • Automated Inventory Counting: Drones equipped with barcode scanners or RFID readers can cruise through the warehouse to identify and record inventory data, significantly reducing manual labor and time. For example, drones are capable of scanning hundreds of locations per hour.
  • Cyclical Stocktaking: Drones can perform regular inventory audits without disrupting day-to-day operations.
  • Enhanced Safety: By accessing high or narrow areas, drones reduce the risks associated with manual handling.
  • Real-time Updates and Management: These systems provide real-time inventory data, especially useful for managing perishable items using the First-Expire-First-Out (FEFO) strategy to improve efficiency [23].
Current evidence shows that drones excel in inspecting elevated warehouse areas. However, challenges remain in terms of cost, safety, regulatory compliance, and compatibility with warehouse layout [22]. Therefore, although drone-assisted inventory systems hold great potential in warehouse automation, their effectiveness depends on overcoming technological, economic, and operational barriers, particularly in developing solutions for scanning stacked inventory.
Despite the promise of drone-assisted inventory systems, their widespread application is hindered by several limitations, summarized in Table 2. Moreover, the integration of Artificial Intelligence (AI) and Machine Learning (ML) has significantly enhanced the accuracy of label scanning, reaching precision levels of 97% to 99% [25], which constitutes an unexpected but noteworthy advantage.
Table 2: Limitations of Drone-Assisted Inventory Systems
Limitation Description
Battery Life Constraints Limited battery life requires efficient path planning or the deployment of multiple drones to cover large areas.
Indoor Navigation The absence of GPS signals indoors necessitates the use of costly SLAM (Simultaneous Localization and Mapping) or beacon systems for navigation.
Stacked Inventory Limitation In densely stored or vertically stacked environments, drones struggle to scan labels obscured by other items, leading to decreased accuracy.
Operational Safety Robust obstacle avoidance mechanisms are essential to ensure safe operation.
High Initial Cost The cost of hardware, software, and integration with Warehouse Management Systems (WMS) may be prohibitive for small businesses.
2.4 State-of-the-Art Comparison
In the realm of supply chain optimization, selecting an effective inventory management system hinges on a systematic evaluation of key criteria and their assigned weights. [29] underscore the importance of inventory accuracy and real-time data, while [30] stress the critical role of information accuracy and timeliness. Drawing from these foundational works, we identified five essential evaluation criteria: cost efficiency, inventory accuracy, implementation complexity, real-time performance, and scalability. The Analytic Hierarchy Process (AHP) was utilized to determine their relative importance through expert pairwise comparisons.
2.4.1 AHP Derivation Process
AHP employs a 1-9 scale to quantify expert judgments, where 1 signifies equal importance and 9 denotes extreme importance. Experts in this study prioritized cost efficiency and inventory accuracy as key drivers in reducing supply chain inefficiencies. The weight is calculated as follows:
  1. 1.
    Matrix Normalization: Each column is normalized by dividing its elements by the column sum. For example, the first column sums to
    $1+\frac{1}{3}+\frac{1}{5}+\frac{1}{4}+\frac{1}{4}=2.033,$
    yielding normalized values such as 1/2.033=0.4921/2.033=0.4921 / 2.033 = 0.492, (1/3)/2.033=0.164(1/3)/2.033=0.164( 1 / 3 ) / 2.033 = 0.164, etc.
  2. 2.
    Eigenvector Calculation: Weights are computed by averaging the rows of the normalized matrix, yielding cost efficiency 47.6%, inventory accuracy 21.0%, implementation complexity 7.1%, real-time performance 12.1%, and scalability 12.1%.
  3. 3.
    Consistency Check: The consistency ratio (CR) of 0.045, below the 0.1 threshold, confirms the matrix’s reliability.
The resulting judgment matrix is shown in Table 3.
Table 3: Judgment Matrix for Evaluation Criteria
Cost Efficiency Inventory Accuracy Implementation Complexity Real-Time Performance Scalability
Cost Efficiency 1 3 5 4 4
Inventory Accuracy 1/3 1 3 2 2
Implementation Complexity 1/5 1/3 1 1/2 1/2
Real-Time Performance 1/4 1/2 2 1 1
Scalability 1/4 1/2 2 1 1
2.4.2 System Evaluation Results
Three inventory management systems were assessed using these weights on a 10-point scale, with results presented in Table 4. The Barcode-Flink-AI system emerged as the top performer, excelling in cost efficiency and inventory accuracy—criteria with the highest weights.
Table 4: System Evaluation Scores
System Weighted Score
Barcode-Flink-AI System 9.15
IoT-Based System 5.6
Drone-Assisted System 4.3
2.4.3 Robustness Analysis Under Weight Variations
A sensitivity analysis was conducted to evaluate the Barcode-Flink-AI system’s adaptability by adjusting weight distributions to reflect varying supply chain priorities, such as raising real-time performance to 30% and lowering cost efficiency to 35% or increasing scalability to 25% while reducing inventory accuracy to 15%. Across these scenarios, the Barcode-Flink-AI system’s score ranged from 8.8 to 9.3, consistently outperforming competitors. This resilience highlights its robustness and versatility.
2.4.4 Comparative Analysis Across Key Dimensions
Table 5 compares the Barcode-Flink-AI system with IoT-based and drone-assisted systems across critical dimensions, emphasizing its advantages in cost, integration, and adaptability.
Table 5: Comparison of Inventory Management Systems Across Key Dimensions
Dimension Barcode-Flink-AI System IoT-Based System Drone-Assisted System
Cost Low ($\sim$18of RFID costs) High (sensors, cloud infrastructure) High (hardware, navigation)
Integration Seamless with WMS/ERP Complex; network-dependent Requires specialized navigation
Real-Time Performance Near real-time (99.2% accuracy) Real-time; network-reliant Real-time; battery-limited
Ease of Use Minimal training required Technical expertise needed Drone skills required
Adaptability Flexible across layouts Suited for large operations Limited to drone-friendly setups
Accuracy Maximum 99.2% with AI detection High with proper setup 97%-99% with AI
@@ 2.5 Research Gap
To fill this gap, our system combines the best of barcode technology with cutting-edge stream processing and AI capabilities, offering a more balanced solution that minimizes costs, reduces errors, and scales efficiently. We compare our system to RFID-based, traditional manual inventory counting, and barcode-based systems in Table 6.
@@
Feature RFID-Based System Manual Stocktaking Proposal (Flutter + Flink + AI)
@@
Stocktaking Speed Fast (affected by signal interference) Slow (Manual item-by-item scanning) Fast (Automated barcode scanning + AI anomaly detection)